# Dynamical Quantum Phase Transition and Thermal Equilibrium in the Lattice Thirring Model

### v1

@@ IV.1 Quantum quench and DQPT
In this work, we investigate DQPT in a quantum quench. A quantum quench is the paradigmatic scenario where to study non-equilibrium dynamics of a quantum system. The system is assumed to be prepared in a certain initial state $|\Psi_{1}\rangle$, typically the ground state of a Hamiltonian $H_{1}$. From time $t=0$ on, it is evolved with a different Hamiltonian $H_{2}$. The evolved state can be written as $|\Psi(t)\rangle=e^{-itH_{2}}|\Psi_{1}\rangle$. The key ingredient in the study of DQPT is the Loschmidt amplitude,
@@
which is analogous to the boundary partition function [63] $Z_{B}(z)=\langle\Psi_{1}|e^{-zH_{2}}|\Psi_{1}\rangle$ in the complex plane $z\in\mathbb{C}$. The non-analytical points of the free energy density $f(z)=-\lim_{N\to\infty}\frac{1}{N}\log Z_{B}(z)$ are believed to indicate phase transitions [64]. The squared modulus of Loschmidt amplitude in Eq. (10), the Loschmidt echo, quantifies the return probability, i.e. the probability of the system being found in the initial state after a certain time $t$. In the thermodynamic limit ($N\rightarrow\infty$), the Loschmidt echo will approach $e^{-Nr(t)}$, with the rate function
@@
This function can be interpreted as a dynamical free energy, and can exhibit non-analyticities. Dynamical quantum phase transitions (DQPT) are defined as the times when $r(t)$ is non-analytical [2]. In some cases, DQPTs are associated with quenches across an equilibrium phase transition [2, 65], but such connection to equilibrium is not always present [33], which motivates the identification of DQPTs as a genuinely non-equilibrium effect. The potential observable signature of DQPTs in local and string quantities has also been a subject of intensive studies. [66, 67] The complex phenomenology of DQPTs has motivated studiees in multiple models, but the phenomenon is yet far from a full theoretical understanding [3, 17, 27, 68, 69].
Figure 2: The return rate function and the full spectrum of the mixed transfer matrix $T(t)$. The quantity $r_{i}$ is defined in Eq. (15), while $r_{0}$ corresponds to the return rate function defined in Eq. (11). The insets display the full spectrum of the mixed transfer matrix $T(t)$ defined in Eq. (14), with the red cross representing the corresponding dominant eigenvalue. The time $t^{*}$ (indicated by the grey dash-dotted line) denotes when the first DQPT occurs. This plot shows how the non-analyticities in the return rate $r_{0}(t)$ correspond to discontinuous jumps in the phase of the dominant eigenvalue of $T(t)$ (red point). We can notice that in (a), with $(m_{1}a,\Delta_{1})=(0.6,0.1)$ and $(m_{2}a,\Delta_{2})=(0.1,-0.9)$, there are no DQPTs. However, in (b), a small increase in $m_{1}a$ to $0.7$, with the same quench Hamiltonian, induces a DQPT at $t^{*}=7.78.$
The transfer matrix formalism allows for a systematic exploration of the character of DQPTs [17, 25, 26, 27, 29]. We focus here on the case where the initial state is a MPS, and consider directly the translationally invariant case, i.e., the initial state is parametrized by a single rank-3 tensor $A$. The Loschmidt amplitude, for a translationally invariant system of $N$ sites, can then be expressed as
@@ IV.2 Investigation of DQPT in the lattice version of the Thirring model
Here we focus on the Hamiltonian Eq. (2) and consider quenches from the ground state $|\Psi_{1}\rangle$ of the model with a given set of parameters, $H_{1}=H(m_{1}a,\Delta_{1})$. The quench Hamiltonian, with which the state is evolved, will be denoted as $H_{2}=H(m_{2}a,\Delta_{2})$. For a fixed $H_{2}$, we study the real-time evolution of different initial states, $|\Psi_{1}\rangle$. The latter are represented by their MPS approximation, obtained using the variational uMPS (VUMPS) algorithm [57], and the real time evolution according to $H_{2}$ can be numerically simulated using a standard tensor network algorithm [7, 71, 11]. In particular, we use time-dependent variational principle (TDVP) algorithm [72, 21] to obtain the time-dependent tensors of the MPS representation, $A(t)$. We obtain the spectrum of the mixed transfer matrix, $T(t)$, using Eq. (13) and explore the appearance of DQPTs for various quench parameters. To investigate how DQPTs occur, we introduce the following quantity,
@@
where $\{\lambda_{i}\}$ is the spectrum of the mixed transfer matrix $T(t)$, satisfying $|\lambda_{0}|\geq|\lambda_{1}|\geq\cdots$. Notice that $r_{0}$ is exactly return rate function defined in Eq. (11). Fig. 2 shows the results for a selection of simulation parameters. The first DQPT occurs when $t=t^{*}$. The insets show snapshots of the spectrum of $T(t)$, which demonstrate that the non-analytical points in the return rate function in Eq. (11) are associated with a discontinuous jump in the phase of the dominant eigenvalue of this mixed transfer matrix [70].
@@
V DQPTs and finite temperature phase diagram
@@ V.1 Energy threshold and the inverse temperature
As discussed in the previous section, the appearance of a DQPT after a certain quench is not determined by the fact that the quench crossed an equilibrium phase boundary. Hence the ground state of $H_{2}$ is not the most adequate equilibrium reference for all initial states. Since the mean energy of the initial state $E_{1}:=\langle\Psi_{1}|H_{2}|\Psi_{1}\rangle$ is typically much higher than the ground state energy of the quench Hamiltonian $H_{2}$, one does not necessarily expect the time evolved state to be close to the ground state of $H_{2}$, even at the level of local observables. Instead, it is more meaningful to compare the local properties to those of the thermal equilibrium state at the same energy density. If the system equilibrates (in the sense of locally resembling an equilibrium ensemble), the thermal state should correspond to the maximal entropy state compatible with the conserved quantities. Our particular model conserves energy and total magnetization, and we are considering initial states in which the latter vanishes. Thus, the relevant equilibrium state is the Gibbs state with the same energy $E_{1}$, restricted to the subspace of vanishing total magnetization; namely (and up to normalization) $\rho(\beta)\propto P_{s_{z}=0}\ e^{-\beta H_{2}}P_{s_{z}=0}$, where $P_{s}=\sum_{n,{S}_{z}|n\rangle=s|n\rangle}|n\rangle\langle n|$ is the projector onto the sector of total magnetization $s$. The inverse temperature $\beta_{1}:=\beta(E_{1})$ corresponding to the state is fixed by the condition
@@
Figure 3: Relation $E(\beta)$ between the mean energy and the inverse temperature in the Gibbs state restricted to the zero magnetization sector for $(m_{2}a,\Delta_{2})$=(0.1,-0.9) (left) and (0.1,0.9) (right). We display the energy densities of several explored quenches using star symbols, and circles indicate the initial states where our simulations reveal the occurrence of at least one DQPT. Additionally, the dashed line denotes the lowest value of energy density ($E_{c}/N$) at which a DQPT was observed (see also Fig. 4). For $(m_{2}a,\Delta_{2})$=(0.1,-0.9) (left), $E_{c}/N=-0.0968$ and the corresponding inverse temperature $\beta_{c}\approx 0.34$; for $(m_{2}a,\Delta_{2})$=(0.1,0.9) (right), $E_{c}/N=-0.4045$ and $\beta_{c}\approx 1.70$.
The determination of $\beta_{1}$ can be achieved using standard TNS methods, exploiting the fact that the thermal equilibrium state for a local Hamiltonian can be approximated efficiently by a matrix product operator (MPO) [73, 74]. More concretely, given a value of the inverse temperature $\beta$, the canonical purification of the thermal equilibrium ensemble restricted to the vanishing magnetization sector can be written as $|\Psi(\beta)\rangle:=e^{-\beta H_{2}/2}\sum_{n,S_{z}|n\rangle=0}|n\rangle_{\mathrm{S}}|n\rangle_{\mathrm{A}}$, where the sum runs over a basis restricted to the sector $S_{z}=0$, and the subscripts, $\mathrm{S}$ and $\mathrm{A}$, refer respectively to the system and ancillary degrees of freedom. This state, also known as thermofield double state, can be approximated by a MPS using standard tensor network algorithms [6, 7]. This is achieved by preparing a maximally-mixed state of the system and the ancilla, expressed as a simple MPS, where each system site is entangled only with one, next neighbor, ancillary site, and evolving it in imaginary time with an approximation of the operator $e^{-\beta H_{2}/2}$. In order to restrict the ensemble to a sector of fixed magnetization, we write the projector onto the sector $S_{z}=0$ as a MPO, and use as initial mixed state, instead of the identity. Additionally, this MPO can also be applied to the intermediate state during the evolution to ensure that truncation errors due to the approximations do not induce undesired magnetization components. For a finite system of size $N$, the projector onto vanishing magnetization has an exact MPO representation, with bond dimension $N/2$. Applying the aforementioned TNS evolution algorithm in imaginary time to the vectorized projector provides a MPS approximation to $|\Psi(\beta)\rangle$ for varying $\beta$, from which the thermal expectation value of a physical observable $O$ can be computed directly as
@@
For concreteness, in the rest of the section we focus on two quenches, corresponding to Hamiltonians in each of the phases (see Fig. 1). Here, we choose $(m_{2}a,\Delta_{2})=(0.1,-0.9)$ in the gapless phase, and $(m_{2}a,\Delta_{2})=(0.1,0.9)$ in the gapped one. For each of them we approximate the thermal equilibrium state in a system of $N=300$ sites, sufficiently large to neglect finite size effects in this analysis. The corresponding $E(\beta)$ curves are shown in Fig. 3.
@@
To verify this observation, we estimated the value of $E_{1}$ for a grid of $(m_{1}a,\Delta_{1})$ values, $0<m_{1}a\leq 1.5$ and $-1\leq\Delta_{1}\leq 1$, and identified the regions of parameters for which the threshold (denoted as $E_{c}$) is surpassed (see Fig. 4). We conducted this study for both quench Hamiltonians mentioned above. For the gapless Hamiltonian, $(m_{2}a,\Delta_{2})=(0.1,-0.9)$, the value of $E_{1}$ increases smoothly with both $m_{1}a$ and $\Delta_{1}$, and the plane of constant $E_{c}/N\approx-0.0968$, corresponding to the lowest value for which a DQPT was observed. As indicated in Fig. 4a, this divides the parameter space in two connected regions. Consistent with this, we find DQPTs for all initial states sampled from the region with higher energy.
Figure 4: Energy densities corresponding to ground states of varying $m_{1}a$ and $\Delta_{1}$, as a function of these parameters, for the quench to $(m_{2}a=0.1,\,\Delta_{2}=-0.9)$ (left) and $(m_{2}a=0.1,\,\Delta_{2}=0.9)$ (right). In each case, the dots indicate where DQPT has been observed. The horizontal plane shows the energy threshold, $E_{c}/N$, which is also marked as a black line on the color bars. On the right panel, the circles indicate the family of quenches from constant $m_{1}a=0.1$.
Figure 5: The DQPTs on the inverse temperature-time plane. The first, second, third, and fourth occurrences of DQPTs are colored blue, orange, green, and red, respectively. Grey indicates occurrences beyond the fourth. The grey dash-dotted line denotes the critical inverse temperature $\beta_{c}$. (a) We simulate the evolution parameters $m_{2}a=0.1$ and $\Delta_{2}=-0.9$ and perform a scan for the initial parameters ($m_{1}=0.1,0.2,\dots,1.5$; $\Delta_{1}=-0.9,-0.8,\dots,0.9$). For example, the dashed line is the case for $m_{1}a=0.7$ and $\Delta_{1}=0.1$, which is also specifically shown in Fig. 2b. (b) The simulation results for the evolution parameters $m_{2}a=0.1$ and $\Delta_{2}=0.9$, scanning for the initial parameters as follows: ($m_{1}=0$; $\Delta_{1}=-0.9,-0.8,\dots,-0.1,0.9$), ($m_{1}=0.1,0.2,\dots,1.5$; $\Delta_{1}=-0.9,-0.8,\dots,0.9$), ($m_{1}=1,2,\dots,9$; $\Delta_{1}=1,1.1,1.3,1.5,\dots,1.9$) and ($m_{1}=10,20,30,40,60,80,100,1000$; $\Delta_{1}=0.9$). The black dashed line and the red diamond symbols represent the case for $m_{1}a=1.3$ and $\Delta_{1}=3$, where DQPT does not occur at the lowest branch near $t\approx 4$.
@@
In contrast, for the quench into the gapped phase, $(m_{2}a,\Delta_{2})=(0.1,0.9)$, the dependence of $E_{1}$ on $(m_{1}a,\Delta_{1})$ is not monotonic. We find two disjoint regions of parameters with energies above the identified threshold $E_{c}/N\approx-0.4045$, lying on opposite corners of the parameter space as can be seen in Fig. 4b. One region with negative $\Delta_{1}$ and small $m_{1}a$ (overlapping but not coinciding with the critical phase, as shown in Fig. 1), and the other region with positive large $\Delta_{1}$ and $m_{1}a$, deep within the gapped regime. We simulate quenches from both regions, as well as intermediate parameters with energy $E_{1}$ below the threshold. For the first high-energy region, i.e. negative $\Delta_{1}$ and small $m_{1}a$, we consistently observe DQPTs for energies above the threshold. The second region, with large positive $\Delta_{1}$ and $m_{1}a$, however, seems to exhibit DQPTs only at much later times than the first. An extreme case is shown in Fig. 2d, for the quench from $m_{1}a=1000$, $\Delta_{1}=0.9$. We observe that as $m_{1}a$ values decrease, the time at which the first DQPT is observed increases. Due to limitations imposed by truncation errors, we can reliably explore times up to $t\lesssim 14$. Within this window, we only observe DQPTs for points in the parameter regime with masses outside our original grid. We presume that longer times would identify DQPTs also in the rest of the region above the threshold.
@@
This exploration of the parameter space indicates that for states with similar effective temperature, but corresponding to very different values of $(m_{1}a,\,\Delta_{1})$, the DQPTs may occur in very different ways. In order to obtain a more complete picture, we plot in Fig. 5, the times at which we observe DQPTs as a function of the effective temperature of the initial state for all the simulated cases. For each of the initial states, we determine $\beta$ from the energy expectation value $\langle\Psi_{1}|H_{2}|\Psi_{1}\rangle$ through Eq. (16). The figure shows not only the first observed non-analyticity (blue dots), but the times of all subsequently observed DQPTs (with colors indicating the number of the DQPT). These figures are reminiscent of the non-analyticities shown by the complex Loschmidt rate, when defined on the complex $t$ plane for a fixed initial state [3]. In such case, the non-analyticities of the rate in the complex-time plane are analogous to Fisher zeros of a boundary partition function, and determine the presence and time of occurrence of DQPTs [50, 51]. In our case, the initial state varies and we do not evolve it in imaginary time, but the results suggest that our effective temperature does play an equivalent role, allowing us to explore such complex plane from purely real-time data. However, we also observe that some initial states are not sensitive to some of the DQPT branches. For instance, if we select $(m_{1}a,\Delta_{1})=(3,1.3)$ and quenches to $(m_{2}a,\Delta_{2})=(0.1,0.9)$ (indicated by the black dashed line and red diamond marks in Fig. 5b), the DQPT does not occur at the lowest branch near $t\approx 4$. The generality of this connection for other models thus needs further investigation.

### v2

@@ IV.1 Quantum quench and DQPT
In this work, we investigate DQPT in a quantum quench. A quantum quench is the paradigmatic scenario where to study nonequilibrium dynamics of a quantum system. The system is assumed to be prepared in a certain initial state $|\Psi_{1}\rangle$, typically the ground state of a Hamiltonian $H_{1}$. From time $t=0$ on, it is evolved with a different Hamiltonian $H_{2}$. The evolved state can be written as $|\Psi(t)\rangle=e^{-itH_{2}}|\Psi_{1}\rangle$. The key ingredient in the study of DQPT is the Loschmidt amplitude,
@@
which is analogous to the boundary partition function [64] $Z_{B}(z)=\langle\Psi_{1}|e^{-zH_{2}}|\Psi_{1}\rangle$ in the complex plane $z\in\mathbb{C}$. The nonanalytical points of the free energy density $f(z)=-\lim_{N\to\infty}\frac{1}{N}\log Z_{B}(z)$ are believed to indicate phase transitions [65]. The squared modulus of Loschmidt amplitude in Eq. (10), the Loschmidt echo, quantifies the return probability, i.e., the probability of the system being found in the initial state after a certain time $t$. In the thermodynamic limit ($N\rightarrow\infty$), the Loschmidt echo will approach $e^{-Nr(t)}$, with the rate function
@@
This function can be interpreted as a dynamical free energy, and can exhibit nonanalyticities. DQPTs are defined as the times when $r(t)$ is nonanalytical [2]. In some cases, DQPTs are associated with quenches across an equilibrium phase transition [2, 66], but such connection to equilibrium is not always present [33], which motivates the identification of DQPTs as a genuinely nonequilibrium effect. The potential observable signature of DQPTs in local and string quantities has also been a subject of intensive studies. [67, 68] The complex phenomenology of DQPTs has motivated studiees in multiple models, but the phenomenon is yet far from a full theoretical understanding [3, 17, 27, 69, 70].
Figure 2: The return rate function and the full spectrum of the mixed transfer matrix $T(t)$. The quantity $r_{i}$ is defined in Eq. (15), while $r_{0}$ corresponds to the return rate function defined in Eq. (11). The insets display the full spectrum of the mixed transfer matrix $T(t)$ defined in Eq. (14), with the red cross representing the corresponding dominant eigenvalue. The time $t^{*}$ (indicated by the gray dash-dotted line) denotes when the first DQPT occurs. This plot shows how the nonanalyticities in the return rate $r_{0}(t)$ correspond to discontinuous jumps in the phase of the dominant eigenvalue of $T(t)$ (red point). We can notice that in (a), with $(m_{1}a,\Delta_{1})=(0.6,0.1)$ and $(m_{2}a,\Delta_{2})=(0.1,-0.9)$, there are no DQPTs. However, in (b), a small increase in $m_{1}a$ to $0.7$, with the same quench Hamiltonian, induces a DQPT at $t^{*}=7.78.$
The transfer matrix formalism allows for a systematic exploration of the character of DQPTs [17, 25, 26, 27, 29]. We focus here on the case where the initial state is a MPS and consider directly the translationally invariant case, i.e., the initial state is parametrized by a single rank-3 tensor $A$. The Loschmidt amplitude, for a translationally invariant system of $N$ sites, can then be expressed as
@@ IV.2 Investigation of DQPT in the lattice version of the Thirring model
Figure 3: The return rate functions (red) and the chiral condensate (blue) for the cases (a) without and (b) with DQPTs. The parameters we choose here are the same as those in Figs. 2(a) and 2(b).
Here we focus on the Hamiltonian Eq. (II.1) and consider quenches from the ground state $|\Psi_{1}\rangle$ of the model with a given set of parameters, $H_{1}=H(m_{1}a,\Delta_{1})$. The quench Hamiltonian, with which the state is evolved, will be denoted as $H_{2}=H(m_{2}a,\Delta_{2})$. For a fixed $H_{2}$, we study the real-time evolution of different initial states, $|\Psi_{1}\rangle$. The latter are represented by their MPS approximation, obtained using the variational uMPS (VUMPS) algorithm [58], and the real-time evolution according to $H_{2}$ can be numerically simulated using a standard tensor network algorithm [7, 72, 11]. In particular, we use time-dependent variational principle (TDVP) algorithm [73, 21] to obtain the time-dependent tensors of the MPS representation, $A(t)$. We obtain the spectrum of the mixed transfer matrix, $T(t)$, using Eq. (13) and explore the appearance of DQPTs for various quench parameters. To investigate how DQPTs occur, we introduce the following quantity:
@@
where $\{\lambda_{i}\}$ is the spectrum of the mixed transfer matrix $T(t)$, satisfying $|\lambda_{0}|\geq|\lambda_{1}|\geq\cdots$. Notice that $r_{0}$ is exactly return rate function defined in Eq. (11). Figure 2 shows the results for a selection of simulation parameters. The first DQPT occurs when $t=t^{*}$. The insets show snapshots of the spectrum of $T(t)$, which demonstrate that the nonanalytical points in the return rate function in Eq. (11) are associated with a discontinuous jump in the phase of the dominant eigenvalue of this mixed transfer matrix [71].
@@
IV.3 DQPT and thermalization
Notice that DQPTs and (pre-)thermalization happen at different timescales. In our setup, DQPTs can occur at relatively short times. Hence, we do not expect that local observables have equilibrated yet. Indeed, we observe that local observables are still varying strongly at the time slices when the DQPTs are observed. To illustrate this, we plot in Fig. 3 the time evolution of the chiral condensate (or the staggered magnetization in spin model language),
${\chi}\sim\frac{1}{N}\left|\sum_{n}{(-1)^{n}\left(S_{n}^{z}+\frac{1}{2}\right)}\right|,$
in both the cases with or without DQPT. The chiral condensate shows strong time variation at the longest time allowed for our simulation, $t=10$, in both cases, indicating DQPT and prethermalization or thermalization phenomena occur at different timescales.
V DQPTs and finite-temperature phase diagram
@@ V.1 Energy threshold and the inverse temperature
Figure 4: Relation $E(\beta)$ between the mean energy and the inverse temperature in the Gibbs state restricted to the zero magnetization sector for $(m_{2}a,\Delta_{2})=(0.1,-0.9)$ (left:critical phase) and $(0.1,0.9)$ (right:gapped phase). We display the energy densities of several explored quenches using star symbols, and circles indicate the initial states where our simulations reveal the occurrence of at least one DQPT. Additionally, the dashed line denotes the lowest value of energy density ($E_{c}/N$) at which a DQPT was observed (see also Fig. 5). For $(m_{2}a,\Delta_{2})=(0.1,-0.9)$ (left), $E_{c}/N=-0.0968$ and the corresponding inverse temperature $\beta_{c}\approx 0.34$; for $(m_{2}a,\Delta_{2})=(0.1,0.9)$ (right), $E_{c}/N=-0.4045$ and $\beta_{c}\approx 1.70$.
As discussed in the previous section, the appearance of a DQPT after a certain quench is not determined by the fact that the quench crossed an equilibrium phase boundary. Hence the ground state of $H_{2}$ is not the most adequate equilibrium reference for all initial states. Since the mean energy of the initial state $E_{1}:=\langle\Psi_{1}|H_{2}|\Psi_{1}\rangle$ is typically much higher than the ground-state energy of the quench Hamiltonian $H_{2}$, one does not necessarily expect the time-evolved state to be close to the ground state of $H_{2}$, even at the level of local observables. Instead, it is more meaningful to compare the local properties to those of the thermal equilibrium state at the same energy density. If the system equilibrates (in the sense of locally resembling an equilibrium ensemble), then the thermal state should correspond to the maximal entropy state compatible with the conserved quantities  [^1] A detailed investigation of the equilibration process is beyond the scope of this work.. Our particular model conserves energy and total magnetization, and we are considering initial states in which the latter vanishes. Thus, the relevant equilibrium state is the Gibbs state with the same energy $E_{1}$, restricted to the subspace of vanishing total magnetization; namely (and up to normalization) $\rho(\beta)\propto P_{s_{z}=0}\ e^{-\beta H_{2}}P_{s_{z}=0}$, where $P_{s}=\sum_{n,{S}_{z}|n\rangle=s|n\rangle}|n\rangle\langle n|$ is the projector onto the sector of total magnetization $s$. The inverse temperature $\beta_{1}:=\beta(E_{1})$ corresponding to the state is fixed by the condition
@@
The determination of $\beta_{1}$ can be achieved using standard TNS methods, exploiting the fact that the thermal equilibrium state for a local Hamiltonian can be approximated efficiently by a matrix product operator (MPO) [75, 76]. More concretely, given a value of the inverse temperature $\beta$, the canonical purification of the thermal equilibrium ensemble restricted to the vanishing magnetization sector can be written as $|\Psi(\beta)\rangle:=e^{-\beta H_{2}/2}\sum_{n,S_{z}|n\rangle=0}|n\rangle_{\mathrm{S}}|n\rangle_{\mathrm{A}}$, where the sum runs over a basis restricted to the sector $S_{z}=0$, and the subscripts $\mathrm{S}$ and $\mathrm{A}$ refer respectively to the system and ancillary degrees of freedom. This state, also known as thermofield double state, can be approximated by a MPS using standard tensor network algorithms [6, 7]. This is achieved by preparing a maximally-mixed state of the system and the ancilla, expressed as a simple MPS, where each system site is entangled only with one, next neighbor, ancillary site, and evolving it in imaginary time with an approximation of the operator $e^{-\beta H_{2}/2}$. In order to restrict the ensemble to a sector of fixed magnetization, we write the projector onto the sector $S_{z}=0$ as a MPO and use as initial mixed state, instead of the identity. Additionally, this MPO can also be applied to the intermediate state during the evolution to ensure that truncation errors due to the approximations do not induce undesired magnetization components. For a finite system of size $N$, the projector onto vanishing magnetization has an exact MPO representation, with bond dimension $N/2$. Applying the aforementioned TNS evolution algorithm in imaginary time to the vectorized projector provides a MPS approximation to $|\Psi(\beta)\rangle$ for varying $\beta$, from which the thermal expectation value of a physical observable $O$ can be computed directly as
@@
For concreteness, in the rest of the section we focus on two quenches, corresponding to Hamiltonians in each of the phases (see Fig. 1). Here we choose $(m_{2}a,\Delta_{2})=(0.1,-0.9)$ in the gapless phase and $(m_{2}a,\Delta_{2})=(0.1,0.9)$ in the gapped one. For each of them we approximate the thermal equilibrium state in a system of $N=300$ sites, sufficiently large to neglect finite-size effects in this analysis. The corresponding $E(\beta)$ curves are shown in Fig. 4.
Figure 5: Energy densities corresponding to ground states of varying $m_{1}a$ and $\Delta_{1}$, as a function of these parameters, for the quench to the critical phase $(m_{2}a=0.1,\,\Delta_{2}=-0.9)$ (left) and the gapped phase $(m_{2}a=0.1,\,\Delta_{2}=0.9)$ (right). In each case, the dots indicate where DQPT has been observed. The horizontal plane shows the energy threshold, $E_{c}/N$, which is also marked as a black line on the color bars. On the right panel, the circles indicate the family of quenches from constant $m_{1}a=0.1$.
@@
To verify this observation, we estimated the value of $E_{1}$ for a grid of $(m_{1}a,\Delta_{1})$ values, $0<m_{1}a\leq 1.5$ and $-1\leq\Delta_{1}\leq 1$, and identified the regions of parameters for which the threshold (denoted as $E_{c}$) is surpassed (see Fig. 5). We conducted this study for both quench Hamiltonians mentioned above. For the gapless Hamiltonian, $(m_{2}a,\Delta_{2})=(0.1,-0.9)$, the value of $E_{1}$ increases smoothly with both $m_{1}a$ and $\Delta_{1}$, and the plane of constant $E_{c}/N\approx-0.0968$, corresponding to the lowest value for which a DQPT was observed. As indicated in Fig. 5(a), this divides the parameter space in two connected regions. Consistent with this, we find DQPTs for all initial states sampled from the region with higher energy.
@@
In contrast, for the quench into the gapped phase, $(m_{2}a,\Delta_{2})=(0.1,0.9)$, the dependence of $E_{1}$ on $(m_{1}a,\Delta_{1})$ is not monotonic. We find two disjoint regions of parameters with energies above the identified threshold $E_{c}/N\approx-0.4045$, lying on opposite corners of the parameter space as can be seen in Fig. 5(b). One region with negative $\Delta_{1}$ and small $m_{1}a$ (overlapping but not coinciding with the critical phase, as shown in Fig. 1), and the other region with positive large $\Delta_{1}$ and $m_{1}a$, deep within the gapped regime. We simulate quenches from both regions, as well as intermediate parameters with energy $E_{1}$ below the threshold. For the first high-energy region, i.e., negative $\Delta_{1}$ and small $m_{1}a$, we consistently observe DQPTs for energies above the threshold. The second region, with large positive $\Delta_{1}$ and $m_{1}a$, however, seems to exhibit DQPTs only at much later times than the first. An extreme case is shown in Fig. 2(d), for the quench from $m_{1}a=1000$, $\Delta_{1}=0.9$, for which we observe a DQPT only close to $t\approx 10$ We observe that as $m_{1}a$ values decrease, the time at which the first DQPT is observed increases even more, so that we need to run long-time simulations to analyze this region. Due to limitations imposed by truncation errors, we can reliably explore times up to only $t\lesssim 14$. Within this window, we only observe DQPTs for points in the parameter regime with masses outside our original grid. We presume that longer times would identify DQPTs also in the rest of the region above the threshold.
Figure 6: The DQPTs on the inverse temperature–time plane. The first, second, third, and fourth occurrences of DQPTs are colored blue, orange, green, and red, respectively. Gray indicates occurrences beyond the fourth. The gray dash-dotted line denotes the critical inverse temperature $\beta_{c}$. (a) We simulate the evolution parameters $m_{2}a=0.1$ and $\Delta_{2}=-0.9$ (critical phase) and perform a scan for the initial parameters ($m_{1}a=0.1,0.2,\dots,1.5$; $\Delta_{1}=-0.9,-0.8,\dots,0.9$). For example, the dashed line is the case for $m_{1}a=0.7$ and $\Delta_{1}=0.1$, which is also specifically shown in Fig. 2(b). (b) The simulation results for the evolution parameters $m_{2}a=0.1$ and $\Delta_{2}=0.9$ (gapped phase), scanning for the initial parameters as follows: ($m_{1}a=0$; $\Delta_{1}=-0.9,-0.8,\dots,-0.1,0.9$), ($m_{1}a=0.1,0.2,\dots,1.5$; $\Delta_{1}=-0.9,-0.8,\dots,0.9$), ($m_{1}a=1,2,\dots,9$; $\Delta_{1}=1,1.1,1.3,1.5,\dots,1.9$), and ($m_{1}a=10,20,30,40,60,80,100,1000$; $\Delta_{1}=0.9$). These large masses are chosen to explore the large $\Delta_{1}$, large $m_{1}a$ corner of Fig. 5(b), while ensuring that DQPT occurs within our simulation window ($t<14$). The black dashed line and the red diamond symbols represent the case for $m_{1}a=1.3$ and $\Delta_{1}=3$, where DQPT does not occur at the lowest branch near $t\approx 4$.
@@
This exploration of the parameter space indicates that for states with similar effective temperature, but corresponding to very different values of $(m_{1}a,\,\Delta_{1})$, the DQPTs may occur in very different ways. In order to obtain a more complete picture, we plot in Fig. 6, the times at which we observe DQPTs as a function of the effective temperature of the initial state for all the simulated cases. For each of the initial states, we determine $\beta$ from the energy expectation value $\langle\Psi_{1}|H_{2}|\Psi_{1}\rangle$ through Eq. (16). The figure shows not only the first observed nonanalyticity (blue dots), but the times of all subsequently observed DQPTs (with colors indicating the number of the DQPT). These figures are reminiscent of the nonanalyticities shown by the complex Loschmidt rate, when defined on the complex $t$ plane for a fixed initial state [3]. In such case, the nonanalyticities of the rate in the complex-time plane are analogous to Fisher zeros of a boundary partition function and determine the presence and time of occurrence of DQPTs [51, 52]. In our case, the initial state varies and we do not evolve it in imaginary time, but the results suggest that our effective temperature does play an equivalent role, allowing us to explore such complex plane from purely real-time data. However, we also observe that some initial states are not sensitive to some of the DQPT branches. For instance, if we select $(m_{1}a,\Delta_{1})=(3,1.3)$ and quenches to $(m_{2}a,\Delta_{2})=(0.1,0.9)$ [indicated by the black dashed line and red diamond marks in Fig. 6(b)], then the DQPT does not occur at the lowest branch near $t\approx 4$. The generality of this connection for other models thus needs further investigation.